(vi) z = TEM (vii) 2 = 22 + 231

Find R and angle. Z1 =8+3i, Z2 =2+3i,
Z3 =9-((2)^1/2 )i.
Answer #1

Given that Z, = 8+31 , 2q = 2+31 . Z3 - 9-N2 i ving Given that z e zkrizi 2-(x+y) : reie Magnitude 8 = 121 = {x?ty? . Angle ovi) Given that z = 121 - (8+3) { Let us solve it alternative method Square on both side. (6+3819 = 2827381-9 flat it; (2246)x: 土望,是 3-73-2 : 73 - 经: { 16- * - 绥 | Z + (+ ) | 12 (15) |??- Mtimes) G-sh) | z. E(n-15)。 (P.Mary 816746 2: 0.027/T - 0.6787

251112 + (3-12) Z= N121 + 9 +2-6N2 Z = N132-6NZ = N121.515 observe that is not a modulus so It is just magnitude. so r = N123